OpenStudy (anonymous):

Solve 5^x+5=9^x help plz im lost

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

5x+5-9x=0 thats yo answer

OpenStudy (anonymous):

no look dont you see the exponent its 5 ^x+5=9 ^x

OpenStudy (anonymous):

5^x+5-9^x=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 5^x+5=9^x

OpenStudy (anonymous):

\[5^x +5 = 9^x\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, take ln both sides, or log , either way is ok, do it,

OpenStudy (anonymous):

its hard

OpenStudy (anonymous):

choose log

OpenStudy (anonymous):

not hard, it is easy

OpenStudy (anonymous):

wait its not that

OpenStudy (anonymous):

its 5^x+5 = 9^x

OpenStudy (anonymous):

\[\log (5^x + 5) = \log 9^x\]

OpenStudy (anonymous):

thats and exponent

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

no you put the wrong question

OpenStudy (anonymous):

the x+5 is together

OpenStudy (anonymous):

ok, it 's easier

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\log 5^(x+5) = \log 9\] ---> (x+5) log 5 = log 9, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, xlog5 +5 log5 = log9, ok?

OpenStudy (anonymous):

cant you rewrite it 9^x=5^x+5

OpenStudy (anonymous):

CANNOT

OpenStudy (anonymous):

continue or not?

OpenStudy (anonymous):

yes continue

OpenStudy (anonymous):

good, so xlog5 = log9 - log 25, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the last step is do something to make it neatly, xlog5 = log 9/25 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok, x = \[x = \frac{ \log \frac{ 5 }{ 9 } }{\log 5 }\]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

so whats the answer in decimal form

OpenStudy (anonymous):

not finish yet, friend,

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x = \log_{5}\frac{ 5 }{ 9 } \] it means log (base 5) of 5/9. that's your answer, If you want more neatly, take out calculator and manipulate on it. me done

OpenStudy (anonymous):

but that didnt help i neeed it in decimal form

OpenStudy (anonymous):

plz do it its the last question ill giv medal

OpenStudy (anonymous):

-0.37

OpenStudy (anonymous):

thats not one of the answer choices

OpenStudy (anonymous):

i don't care about the medal, just care whether you can get the stuff or not

OpenStudy (anonymous):

a.-3.63 b.-0.28 c.0.07 d.13.69

OpenStudy (anonymous):

@Luis_Rivera take it over, please, don't just laugh at me like that

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

ok, I messed up somewhere, i

OpenStudy (anonymous):

my bad, sorry,

OpenStudy (anonymous):

x = (log 9/25 )/ log 5

OpenStudy (anonymous):

in decimal form

OpenStudy (anonymous):

0.63

OpenStudy (anonymous):

-0.63

OpenStudy (anonymous):

not one of the choices

OpenStudy (anonymous):

i think you read the question wrong

OpenStudy (anonymous):

ok, I'm with you now, repost the question in new post, I'll be there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5^x+5=9^x

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

I try to guide you on the simplest way.

OpenStudy (anonymous):

ok continue

OpenStudy (anonymous):

ok, take log both side, you have log 5^(x+5) = log9^x

OpenStudy (anonymous):

yes thats the right problem

OpenStudy (anonymous):

take exponent down, you have (x+5) log 5 = xlog 9

OpenStudy (anonymous):

xlog5 + 5log5 -xlog9 =0, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

xlog5 -xlog9 = -5log5, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

factor x out x(log5 -log9) = -5log5 ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so x = \[x = \frac{ -5\log5 }{ \log5 -\log9 }\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ok decimal form

OpenStudy (anonymous):

you take out the calculator and manipulate, why I have to do it for you?

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

13.69

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

im going to bookmark it ty man

OpenStudy (anonymous):

ha, !!! you really drive me crazy

OpenStudy (anonymous):

I will avoid you next time,

OpenStudy (anonymous):

joking

OpenStudy (anonymous):

ok im in high school wbu